OpenStudy (anonymous):

how do i divide (-4/5) by (-1.6) i really need help

OpenStudy (anonymous):

-4/5=-0.8 now divide it by -1.6

Parth (parthkohli):

\[ \color{Black}{\Rightarrow {-4 \over 5} \div {-16 \over 10}}\]

Parth (parthkohli):

@theyatin He has mentioned -1.6.. hmm..

OpenStudy (anonymous):

How do i convert -4/5 into a decimal?

OpenStudy (anonymous):

@ParthKohli so did i

Parth (parthkohli):

No need to. \[\color{Black}{\Rightarrow {-4 \over 5} \div {-16 \over 10} \Longrightarrow {-4 \over 5} \times{10 \over -16}}\]

OpenStudy (anonymous):

Thsnk You but how do i convert the decimal into a fraction?

Parth (parthkohli):

\[ \color{Black}{\Rightarrow {\cancel{-4} \over \cancel{5}} \times {\cancel{10}^2 \over \cancel{-16^4}}}\]

Parth (parthkohli):

@MisaelJnr Please Google that :)

OpenStudy (anonymous):

So, then the answer is 40/80 simplified into 2?

Parth (parthkohli):

Hmm. You just got a little mistake in simplifying that. You simplified \(\Large {80 \over 40}\) instead of \(\Large {40 \over 80}\).

OpenStudy (anonymous):

So the answer is 5/10